# Aligning to Thousands of Preferences via System Message Generalization

### v1

@@ 5 Experimental results
5.1 Can Janus generate more personalized responses effectively?
@@
Figure 3 illustrates the win rates of model responses on Multifaceted Bench, reflecting human judgments on their ability to handle unseen preferences. Janus 7B outperforms Mistral Instruct 7B v0.2 in 36.8% of comparisons and is favorably rated in 74.8% of instances. In contrast, Mistral Instruct 7B v0.2 surpasses Janus 7B in 12.4% of cases. When compared with proprietary models optimized for preferences, Janus 7B exceeds GPT-3.5-Turbo-0125 in 47% of cases and is outperformed in 19.4% of them. Against GPT-4-0613, Janus 7B leads in 37.5% of comparisons, while GPT-4-0613 excels in 27.3% of them. Overall, Janus matches or surpasses the performance of Mistral 7B Instruct v0.2, GPT-3.5 Turbo, and GPT-4, 75.2%, 72.4%, and 66.4% of the time, respectively. These results underscore Janus’s effectiveness in generating contextually personalized responses that adapt well to unseen preferences, demonstrating its scalability and applicability in diverse real-world settings.
Using the LLM-as-a-Judge method, our Janus suite consistently surpasses other models in multifacetedness on the Multifaceted Bench. As shown in Table 2, Janus 7B scores 38.8% higher than the Mistral 7B v0.2 base model and 4.8% above its instruction-tuned counterpart, Mistral 7B Instruct v0.2. It also performs competitively with larger instruction-tuned models, including LLaMA 2 Chat 70B, Mixtral 8x7B Instruct v0.1, and LLaMA 3 Instruct 70B, only slightly trailing the latter. Against LLaMA 2-Chat-70B and Mixtral 8x7B Instruct v0.1, Janus registers average score improvements of 7.6% and 2.8%, respectively. Additionally, incorporating preference optimization methods such as DPO with Multifaceted Collection further enhances performance by 1.7%, marking the highest achievement within the Janus suite. When compared with proprietary models designed for human preferences, Janus 7B remains competitive, outscoring GPT-3.5-Turbo and GPT-4 by 5.2% and 1.4%, respectively, although GPT-4-Turbo, which is used in generating Multifaceted Collection, does surpass Janus.
We also aim to assess whether responses by Janus are more diverse compared to other models. We evaluate the diversity of responses from three models on Multifaceted Bench by calculating the average ROUGE-L score among responses for different system messages. As shown in Figure 9, Janus shows the lowest mean and maximum ROUGE-L scores, indicating its ability to generate distinctly varied responses to different prompts, thus displaying greater diversity.
These results confirm that Janus adeptly adjusts to a wide array of preferences by training on system messages tailored to multifaceted and personalized responses. This adaptability, combined with its compatibility with standard preference optimization techniques, highlights Janus’s practical utility and robustness without relying on supplementary models like reward models.
@@
Table 3: Helpfulness benchmarks results. indicates the results from the official leaderboard or paper [12]. Unk. refers to models whose parameter counts are not publicly disclosed. Details regarding the score metric can be found in Section F.3.
@@
5.2 Can Janus also align with generally helpful values?
To assess whether Janus, trained for individual preferences, also meets general helpfulness standards required by current LLM benchmarks, we evaluated it across three benchmarks: AlpacaEval 2.0, MT-Bench, and Arena Hard Auto v0.1. According to Table 3, Janus 7B shows higher LC win rates than similarly sized models in AlpacaEval 2.0, achieving improvements of 9.8%p over Mistral 7B Instruct v0.2, 16.5%p over Gemma 7B Instruct, and 4%p over LLaMA 3 8B Instruct. It also surpasses larger models like Vicuna 33B v1.3 and Mixtral 8x7B Instruct v0.1, as well as the proprietary model GPT-3.5-Turbo, with gains of 10.3%p, 3.2%p, and 4.2%p, respectively. In MT-Bench’s absolute evaluation setting, Janus 7B matches or exceeds other models under 30B in size, scoring 7.7, a noteworthy performance for its model size. Although larger and proprietary models outperform Janus 7B, its results are commendable given its scale. Additionally, in Arena Hard Auto v0.1, Janus 7B scores 20.9, outdoing both smaller LLMs and larger models, including GPT-3.5-Turbo-0125, Vicuna 33B v1.3, and Tulu 2 + DPO 70B.
These results indicate that Janus 7B excels across all helpfulness benchmarks relative to other models. Notably, despite being trained specifically to generate personalized responses when provided context through system messages, Janus does not falter in general helpfulness. In fact, its performance suggests that system message generalization not only supports the creation of personalized LLMs but also acts as an effective novel method for improving alignment with what humans generally perceives as helpful. This dual success in multifaceted and helpfulness benchmarks demonstrates that Janus’s approach to training could redefine how LLMs align with user expectations.
5.3 Can Janus maintain diversity while also ensuring safety?
Janus is designed to reflect a broad range of individual values while prioritizing safety by maintaining consistent standards for harmlessness. We evaluate the safety of Janus by analyzing its toxicity using metrics from RealToxicityPrompts [14]. As detailed in Table 4, Janus 7B shows an average maximum toxicity of 0.256 and a toxicity probability of 0.061, significantly lower than other instruction-tuned models. This not only confirms its lower toxicity but also shows that unlike traditional method [42] where reducing toxicity could compromise performance, Janus 7B manages to maintain lower toxicity while still achieving high fluency and diversity scores. These findings underscore Janus’s effectiveness in balancing diversity, helpfulness, and safety, making it a robust and versatile model.
Table 4: Harmlessness analysis on RealToxicityPrompts. indicates the results from previous work [42]. Details regarding the score metric can be found in Section F.4.
@@ 5.3 Can Janus maintain diversity while also ensuring safety?
GPT-2 [51] 0.527 0.520 11.31 0.85 0.85
PPLM [11] 0.520 0.518 32.58 0.86 0.86
GeDi [31] 0.363 0.217 60.03 0.84 0.83
DExperts [40] 0.314 0.128 32.41 0.84 0.84
DAPT [16] 0.428 0.36 31.21 0.84 0.84
PPO [63] 0.218 0.044 14.27 0.8 0.84
Quark [42] 0.196 0.035 12.47 0.8 0.84
Mistral 7B Instruct v0.2 0.292 0.113 19.43 0.92 0.92
LLaMA 3 Instruct 8B 0.296 0.118 28.88 0.92 0.92
Janus 7B 0.256 0.061 14.58 0.93 0.95
@@ 6 Analysis
We conduct several ablation studies and analyses to closely examine the behavior of Janus. In Section 6.1, we demonstrate that Janus exhibits system prompt generalization compared to other models. Section 6.2 illustrates the benefits of incorporating multifacetedness into the system message. In Section 6.3, we show that Janus can be a reward model aligned with diverse values. Furthermore, in Section G.1, we conduct a qualitative analysis of actual response examples generated by Janus and other baseline models.
6.1 Janus demonstrates robust performance with or without system messages
@@
We aim to assess how models perform with and without a system message at test time. We provide either no message or a multifaceted one. Despite training with a multifaceted message, our model shows only a minor performance drop when no message is given, as seen in Table 8. Similarly, a proprietary LLM, although not specifically trained with these messages, experiences a slight decline in performance yet still scores highest under both conditions. Conversely, models like Mistral 7B Instruct v0.2 and Mistral 7B v0.2 show significant capability with a multifaceted message but face substantial decreases in performance without it. This indicates that Janus inherently possesses multifaceted capabilities and aligns well with human values, suggesting it can generate quality responses regardless of the presence of a system message. Moreover, the superior performance of the proprietary model confirms the effectiveness of using synthetic data, referred to as Multifaceted Collection, in creating personalized LLMs.
6.2 Janus benefits from incorporating multifaceted system messages
@@
We conducted an ablation study to explore how multifacetedness in training data affects model performance, training our model in four distinct manners: 1) without a system message to generate generally helpful responses, 2) with a default system message to elicit helpful responses, 3) without a system message to produce multifaceted responses, and 4) using both a system message and multifaceted responses from Multifaceted Collection. List of the default system messages we used can be seen in Table 9. We evaluated these models using Multifaceted Bench and MT-Bench. As detailed in Table 10, training with a multifaceted system message enhances performance in both multifacetedness and helpfulness. Models with a default system message improve in helpfulness but show a slight reduction in multifacetedness. Training to generate multifaceted responses improves both metrics, though it underperforms compared to models trained with an explicit system message. This suggests that training without a system message, which demands generating personalized responses without specific prompts, poses challenges in effectively learning the nuances of human preferences. We refer to the system messages used by existing Large Language Models (LLMs) as ‘default system messages’.
6.3 Janus can be an efficient personalized reward model
In this study, we explore the effectiveness of training a reward model with Multifaceted Collection to enhance model performance on Multifaceted Bench. We initially train Janus* 7B using only 65k samples from Multifaceted Collection and then train the reward model, Janus-RM 7B, with additional data. The effects of these training approaches are evaluated using best-of-n sampling, with details available in section E.2. In Table 11, applying best-of-64 sampling to Janus* 7B increases the average score from 4.14 to 4.28. When used with Janus+DPO 7B, the score further improves from 4.24 to 4.31, marking the highest performance within the Janus suite. This improvement suggests that Janus-RM 7B effectively aligns with multifaceted preferences and can provide tailored rewards to meet individual user requests, enhancing personalized response generation without the need for training multiple models on diverse values as previously done.
7 Conclusion
In this paper, we introduce a novel method to align LLMs with diverse user preferences without requiring continual retraining for each individual preference. Our approach utilizes a unique system message protocol that guides LLMs to produce responses tailored to specific, nuanced user preferences. To accomplish this, we also create a comprehensive training dataset, named Multifaceted Collection, which is enriched with a variety of system messages to enhance the model’s adaptability to different user instructions. The personalized LLM trained with this dataset, Janus 7B, significantly surpasses baseline models in various benchmarks, showcasing its effectiveness in adapting to diverse, unseen preferences. Notably, it also demonstrates superior performance compared to similar-sized baseline models in benchmarks focused on helpfulness. Through various analyses, we establish that Janus not only retains its capability for generalization across system messages and diversity but also maintains harmlessness. We believe our work contributes to developing AI systems that respect the values of both the majority and individual user preferences.
Appendix A Details of Multifaceted Collection construction
A.1 Instruction sampling and filtering

### v2

@@ 5 Experimental results
5.1 Unseen multifaceted preference
@@
Humans prefer Janus for personalized responses.
Figure 3 illustrates the win rates of model responses on Multifaceted Bench, reflecting human judgments on their ability to handle unseen preferences. Janus achieves tie+win rate [^6] We only include tie scenarios where both model responses are good. of 75.2%, 72.4%, and 66.4% against Mistral 7B Instruct v0.2, GPT-3.5-Turbo-0125, and GPT-4-0613, respectively.
Janus consistently surpasses other models when judged by an LLM.
Table 2 shows LLM-as-a-Judge evaluation results on Multifaceted Bench. Janus 7B scores higher than its base model, Mistral 7B v0.2 (+33%), and its instruction-tuned counterpart, Mistral 7B Instruct v0 (+4.6%). It also competes well with larger models like LLaMA 2 Chat 70B (+9%) and Mixtral 8x7B Instruct v0.1 (+3.8%), only slightly trailing behind LLaMA 3 Instruct 70B. Janus remain competitive against proprietary models like GPT-3.5-Turbo (+7.6%) and GPT-4 (+3.6%). Incorporating preference optimization methods like DPO with Multifaceted Collection further enhances performance by 0.4%, marking the highest achievement within the Janus suite.
Reward modeling on Multifaceted Collection can effectively distinguish between different values in responses.
We study how well our reward model, Janus-RM 7B, can additionally improve generation of Janus models through Best-of-64 sampling. As seen in Table 3, applying Best-of-64 sampling to Janus* 7B increases the average score from 4.14 to 4.28. When used with Janus+DPO 7B, the score further improves from 4.24 to 4.31, marking the highest performance within the Janus suite. The improvements suggests that a single reward model trained on diverse values can enhance multifaceted response generation without the need of multiple reward models.
Overall, Janus adeptly adjusts to a wide array of preferences by training on system messages tailored to multifaceted responses. This adaptability, combined with its compatibility with standard preference optimization techniques, highlights Janus’s practical utility to align to unseen values of individuals.
@@
Janus* 7B 4.41 4.02 4.36 4.08 4.03 4.14
+ Best-of-64 4.41 4.26 4.42 4.17 4.21 4.28
Janus+DPO 7B 4.45 4.13 4.43 4.21 4.17 4.24
+ Best-of-64 4.49 4.24 4.48 4.20 4.28 4.31
Table 4: Helpfulness benchmarks results. indicates the results from the official leaderboard or paper [12]. Unk. refers to models whose parameter counts are not publicly disclosed. Details regarding the score metric can be found in Appendix L.2.
@@
5.2 General helpfulness
Janus outperforms similarly-sized or even larger models in general benchmarks.
Table 4 reports strong performance of Janus across all general helpfulness benchmarks. In AlpacaEval 2.0, Janus 7B shows higher length-controlled (LC) win rates than similarly sized models, Mistral 7B Instruct v0.2 (+9.8%p), Gemma 7B Instruct (+16.5%p), and LLaMA 3 8B Instruct (+4%p). It also surpasses larger models like Vicuna 33B v1.3 (+10.3%p), Mixtral 8x7B Instruct v0.1 (+3.2%p,), as well as the proprietary model GPT-3.5-Turbo (+4.2%p). In MT-Bench’s absolute evaluation setting, Janus 7B matches or exceeds other models under 30B in size, scoring 7.7. Although larger and proprietary models outperform Janus 7B, its results are commendable given its scale. Performance focused on multi-turn scenarios is reported in Appendix A. Additionally in Arena Hard Auto v0.1, Janus 7B scores 20.9, outdoing both smaller LLMs and larger models, including GPT-3.5-Turbo-0125, Vicuna 33B v1.3, and Tulu 2+DPO 70B.
Despite being trained specifically to generate personalized responses when provided preference context through system messages, Janus does not falter, but rather improve in general helpfulness. It is suggested that system message generalization not only supports the creation of personalized LLMs but also improves alignment with what humans generally perceives as helpful.
Table 5: Harmlessness analysis on RealToxicityPrompts. indicates the results from previous work [48]. Details regarding the score metric can be found in Appendix L.3.
@@ 5.2 General helpfulness
GPT-2 [60] 0.53 0.52 11.31 0.85 0.85
PPLM [11] 0.52 0.52 32.58 0.86 0.86
GeDi [34] 0.36 0.22 60.03 0.84 0.83
DExperts [44] 0.31 0.12 32.41 0.84 0.84
DAPT [18] 0.43 0.36 31.21 0.84 0.84
PPO [75] 0.22 0.04 14.27 0.80 0.84
Quark [48] 0.12 0.04 12.47 0.80 0.84
Mistral 7B Instruct v0.2 0.29 0.11 19.43 0.92 0.92
LLaMA 3 Instruct 8B 0.30 0.12 28.88 0.92 0.92
Janus 7B 0.26 0.06 14.58 0.93 0.95
5.3 Diversity and harmlessness
Janus simultaneously achieves low toxicity and high diversity.
Table 5 shows RealToxicityPrompts [16] results of decoding-time algorithms and instruction-tuned models. Janus 7B shows a significantly lower average maximum toxicity and toxicity probability than other instruction-tuned models. Moreover, unlike traditional methods such as [48] where reducing toxicity could compromise performance, Janus 7B manages to maintain lower toxicity while still achieving high fluency and diversity scores. We also observe moderate performance in social bias benchmarks in Appendix B.
@@ 6 Analysis
Figure 4: Length distribution of responses generated by LLMs and reference answer on Multifaceted Bench.
Table 6: Comparison of average and max ROUGE-L scores of three different personalized responses per instruction generated by LLMs in Multifaceted Bench.
Model Avg Max
Mistral 7B Instruct v0.2 0.26 0.31
GPT-4 Turbo 0.28 0.34
Janus 7B 0.23 0.28
Figure 5: Test-time system message ablation results on Multifaceted Bench.
Table 7: Ablating components of multifacetedness in our training data. MF denotes LLM-as-a-Judge scores on Multifaceted Bench. ‘-’ indicates that the data was not used for training.
Base Model System Message Response Average
MF MT-Bench
Mistral 7B Instruct v0.2 - - 3.93 7.23
GPT-4-0613 - - 4.27 9.20
Janus 7B - helpful 3.88 7.41
default helpful 3.87 7.53
- multifaceted 4.01 7.61
multifaceted multifaceted 4.17 7.74
Verbosity of personalized responses
We compare the length distribution of responses from Janus, baseline models, and reference answers on Multifaceted Bench. As illustrated in Figure 4, responses from Janus are, on average, longer than those from Mistral 7B Instruct v0.2, GPT-3.5 Turbo-0125, and GPT-4-Turbo-0125. Since the length distribution of responses from Janus 7B closely matches that of the reference answers, the verbosity can be seen as the result of supervised learning. Still, Table 4 shows that Janus outperforms Mistral 7B Instruct v0.2 and GPT 3.5 Turbo on AlpacaEval 2.0 under the metric that debiases the effect of verbosity.
Diversity of personalized responses
We measure the diversity of responses from three models on Multifaceted Bench by calculating the average textual similarity via ROUGE-L score [43] between responses for three different system messages given one instruction. In Table 6, Janus shows the lowest textual similarity between responses with distinct contexts. This empirically displays the ability to generate varying responses to different prompts as expected.
Effect of excluding multifaceted system messages at test time
@@
We test the robustness of the performance of Janus by providing no multifaceted system messages when generating responses to Multifaceted Bench instructions. Figure 5 shows that Janus shows only a minor performance drop when no message is given, despite being trained only with prompts including system messages. Interestingly, a similar trend is observed across other models that are not even trained with multifaceted system messages, but with varying magnitudes of performance drops. A proprietary LLM from which the ability of Janus was distilled, GPT-4-Turbo-0125, experiences a slight decline in performance yet still scores the highest under both conditions. Conversely, models like Mistral 7B Instruct v0.2 and Mistral 7B v0.2 face substantial decreases. The results imply that Janus can generate quality responses regardless of the presence of a system message. Moreover, the superior performance of the proprietary model confirms the effectiveness of using our synthetic dataset, Multifaceted Collection, in creating personalized LLMs.
Benefit of incorporating multifacetedness during training
@@
We explore how the multifacetedness of system messages and responses affects model performance through four training configurations: 1) no system message with generally helpful responses, 2) default system message with helpful responses, 3) no system message with multifaceted responses, and 4) multifaceted system message with multifaceted responses from Multifaceted Collection. Details of default system messages and response collection are in Appendix F. Table 7 shows results using Multifaceted Bench and MT-Bench, with representative baselines for comparison. Training with multifaceted system messages improves both multifacetedness and helpfulness metrics, while default system messages only enhance helpfulness at the cost of reduced multifacetedness. While generating multifaceted responses without system messages shows some improvement, it underperforms compared to models with explicit system messages. This suggests that generating personalized responses without supervision of verbalized preferences is challenging. Overall, exposing user values in the system message and learning to generate personalized responses based on it is an effective strategy for learning the nuances of human preferences.
We also analyze that Janus captures diverse opinions of the human population and further study the effect of our hierarchical data generation strategy and data/model scaling. Additional analyses are detailed in Appendix C, including qualitative analysis of example model responses in Appendix N.
7 Discussion
We have introduced an approach that utilizes a unique system message protocol to guide language models to generate responses tailored to nuanced user preferences. Multifaceted Collection, an instruction dataset enriched with a variety of system messages, is designed to encompass diverse user values determining preferences. We demonstrate our trained Janus models’ adaptability to unseen multifaceted preferences as well as excellence in general helpfulness and harmlessness. Through analyses, we establish the effectiveness of system message generalization, contributing to developing systems that respect both majority and individual user preferences without requiring per-user fine-tuning.
Broader impact of generalizing system messages instead of user messages
We identify significant potential in utilizing meta-instructions in system messages rather than in user messages to reach many alignment targets. While some meta-instructions are dependent on specific user messages, it can also be more general. For example, the system message illustrated in Figure 2 is applicable to any kind of coding problems. In practical applications, preferences can be programmatically set as meta-instructions when relevant user queries are submitted, offering two key advantages: 1) automatic reflection of common user needs across various instructions, and 2) reduced user burden of specifying a meta-instruction every time. Applications like ChatGPT already offers features like custom instructions777openai.com/index/custom-instructions-for-chatgpt/ and memory controls888openai.com/index/memory-and-new-controls-for-chatgpt/: a user can specify their preferences in the custom instruction by oneself (e.g., When I ask you for math problems, please incorporate visual aids.) or ChatGPT will memorize the details itself from the conversations. We believe that including implicit user preferences in the system message without user specification is an important milestone for seamless chat experience, and a model trained for system message generalization will be necessary.
Limitations and future work
Our work is focused on representing user values that constitute a preference in the form of meta-instructions and training models to follow meta-instructions in the form of system messages. While our approach shows promise, several limitations warrant consideration. The reliance on synthetic training data presents inherent challenges, as direct human evaluation of the entire Multifaceted Collection dataset remains unfeasible. Although the performance of Janus suggests generally high data quality and demonstrates lower toxicity levels compared to baseline benchmarks, the flexibility in preference specification may increase vulnerability to malicious attacks such as jailbreaking. To address these concerns, we incorporate safety considerations by including a harmlessness dimension in every system message and implement safeguards in our data generation process. However, the challenge of capturing implicit preferences from previous conversations to form system messages appropriately and managing system message controls remains application-dependent and beyond our current scope. Future work should explore these challenges through dialogue summarization techniques [95, 91], along with enhancing safety through increased safety-related training data and integration with moderation techniques such as LLaMA-Guard [24].
Acknowledgements
We would like to thank Yongrae Jo, Geewook Kim, Hyeonbin Hwang, and Hoyeon Chang for constructive feedback on our initial draft. This work was partly supported by the LG AI Research grant (Self-improving logical reasoning capabilities of LLMs, 2024, 50%) and the Institute for Information & communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (RS-2024-00398115, Research on the reliability and coherence of outcomes produced by Generative AI, 50%).
Appendix A Additional results on helpfulness
Janus is consistently performant in multi-turn scenarios.
Table 8: Multi-turn scenarios results in MT-Bench.
Model Score
Mistral 7B Instruct v0.2 6.0
LLaMA 3 8B Instruct 6.6
Janus 7B 6.8
To evaluate whether Janus effectively adheres to diverse preferences in multi-turn scenarios, we focus on the multi-turn questions from the MT-Bench, excluding single-turn questions. Table 8 show sthat Janus consistently outperforms the baselines, Mistral 7B Instruct v0.2 and LLaMA-3 8B Instruct, with scores of 6.8 compared to 6 and 6.6, respectively. This aligns with the trends observed in Section 5.2, indicating that Janus remains highly competitive in multi-turn settings.
Appendix B Additional results on harmlessness
Table 9: Results on social bias benchmarks. Details regarding the score metric can be found in Appendix L.3
Model Winogender CrowS-Pairs BBQ (Ambig) BBQ (DisAmbig)
Acc ↑ Likelihood Diff ↓ % Stereotype ↓ Acc ↑ Bias score ↓ Acc ↑ Bias score ↓
Mistral 7B Instruct v0.2 0.61 4.45 67.74 0.11 11.63 0.87 2.52
LLaMA 3 8B Instruct 0.64 4.05 64.52 0.08 12.99 0.88 1.98
Gemma 2 9B IT 0.68 5.44 62.43 0.42 7.62 0.86 1.33
Janus 7B 0.64 4.02 67.68 0.08 12.26 0.86 3.25
Figure 6: Winogender accuracy comparison across models. The ‘gotcha’ scenarios refer to a subset of the dataset where the gender of the pronoun referring to an occupation does not match U.S. statistics on that occupation’s majority gender, i.e., they challenge the model’s reliance on stereotypes more.
Janus does not show critical issues of social bias.
According to the results in Table 9, Janus 7B shows a degree of bias similar to that of LLaMA 3 8B Instruct across several tasks, and is better than Mistral 7B Instruct v0.2 except in BBQ. Figure 6 shows a consistent trend of across models and Janus shows competitive performance. We suggest that there is no striking issues of bias in Janus compared to other models.
Appendix C Additional analyses
C.1 Representativeness of the human population
We analyze if Janus’s distribution over answers is well-calibrated to the human population. We adopt Sorensen et al. [73]’s experiments to test if Janus exhibits less similarity to human populations compared to its base pre-trained model (Mistral 7B v0.2) and its post-aligned counterpart (Mistral 7B Instruct v0.2). Following Sorensen et al. [73], models are evaluated on multiple choice datasets of GlobalOpinionQA [14] and Machine Personality Inventory (MPI) [28], and then we calculate the Jensen-Shannon distance between the human (US and Japan) and model distributions in answer choices, averaged over 5 runs per prompt.
Figure 7: Jensen-Shannon distance between the human (US and Japan) and model distributions in GlobalOpinionQA and MPI answer choices. The probability of the model’s next token prediction (logit) for each answer choice selection is calculated, and the probabilities are averaged over 5 runs of the same 3-shot prompt but with randomized answer choices in the few-shot examples. Janus diverges less from the pre-trained distribution than Mistral 7B Instruct v0.2 does.
(a) GlobalOpinionQA
(b) MPI
Figure 8: Distribution of entropy scores across datasets for each model. The process of computing the probability distribution is explained in Figure 1. Models are either pre-aligned or post-aligned in the same model family; Mistral 7B v0.2 is denoted as ‘Pre’, Mistral 7B Instruct v0.2 as ‘Post’, and Janus 7B as ‘Post-Ours’. Janus is closer to the human and pre-trained distribution while Mistral 7B Instruct v0.2 significantly diverges.
Figure 7 shows the distance between human and model distributions. Echoing the results of [73], aligned models including Janus become more distant to the human population after fine-tuning. Still, Janus exhibits a 2x smaller drop in similarity compared to Mistral 7B Instruct v0.2. We also measure the entropy in Figure 8. This visualizes that Janus is significantly closer to pre-trained and human distributions than Mistral 7B Instruct v0.2 does. These experiments further suggest that our training method can facilitate calibration to diverse individuals.
C.2 Additional ablation studies
Effectiveness of hierarchical data generation strategy
We prompt GPT-4-Turbo-0125 to generate preferences freely (a preference set of four values or a single detailed preference description) and qualitatively compare them to our original hierarchically generated ones. On ten samples, we observe that free-form preference generation can create more topic-specific preferences, but oftentimes they deviate from what we expect preferences to be. Specifically, some generations include preferences irrelevant to the goal of the user instruction (e.g., a preference for nutritional benefits in a math problem illustrated with apples) or attempt to resolve the user request and hint solutions (e.g., explicitly instructing correct implementations for a coding problem).
These problems arise because the model needs to understand and elicit preferences from the human user side, not the assistant side. Our hierarchical data generation strategy allows sufficient control over synthesizing what users would expect in the response. We have decided that preferences for any response would differ under the style, background knowledge, informativeness, and harmlessness dimensions based on various existing literature detailed in Section 3.2. Our method of providing the dimensions in context, coupled with manually crafted seed examples, is instrumental in obtaining high-quality, individualized preferences.
Table 10: Ablation for number of system messages.
# system message per instruction # total train instances mf-AlpacaEval mf-FLASK mf-Koala mf-MT-Bench mf-Self-Instruct
1 66k 4.41 4.01 4.36 4.08 4.03
2 132k 4.42 4.05 4.39 4.1 4.01
3 197k ($\rightarrow$ Janus) 4.05 4.39 4.1 4.17 4.24
Table 11: Base model ablation.
Base pre-trained model mf-AlpacaEval mf-FLASK mf-Koala mf-MT-Bench mf-Self-Instruct
Mistral 7B v0.2 ($\rightarrow$ Janus) 4.43 4.06 4.41 4.11 4.01
LLaMA 2 7B 4.41 4.01 4.41 4.08 4.03
LLaMA 2 13B 4.5 4.3 4.5 4.23 4.08
LLaMA 3 8B 4.5 4.4 4.34 4.31 4.14
Effect of data scaling
We explore how the size of Multifaceted Collection affects performance. Janus 7B is trained on a dataset containing 197k instances, where each user prompt is paired with three different systems and their responses. To evaluate the impact of scaling, we reduce the number of systems and responses per user prompt to two and one, resulting in datasets with 132k and 66k instances, respectively, and train separate models for each. The results across five benchmarks in Table 10 show that the scaling effect is evident: increasing the training data volume leads to improved scores.
Effect of model scaling
We investigate how variations in the type and size of the base model affect performance. Initially, we use Mistral 7B v0.2, but we also experiment with LLaMA models. Table 11 results show that LLaMA 2 7B performs similarly to or slightly worse than Mistral 7B v0.2. However, increasing the model size from 7B to 13B (LLaMA 2 7B vs. 13B) leads to a noticeable performance improvement. Notably, the latest model, LLaMA-3 8B, outperforms the larger LLaMA 2 13B in benchmark scores (LLaMA 3 8B vs. LLaMA 2 13B). These findings indicate that both model size and the capabilities of the base pre-trained model significantly influence performance when applying our method.
C.3 Response quality
Janus 7B generates qualitatively superior responses compared to other models. For example, as shown in Table LABEL:tab:example_alternate_expertise, when given system message that require a straightforward, logic-based approach, Janus 7B accurately follows this directive and constructs logical responses. In contrast, while the responses from GPT-4 are not entirely incorrect, they still fail to adhere closely to the user’s preferences. These results demonstrate that Janus not only performs well on benchmarks but also effectively reflects individual human preferences in the quality of its generated responses. For more examples, please see Appendix N.
Appendix D Details of Multifaceted Collection construction
Instruction sampling and filtering